OpenStudy (anonymous):

In 2007, Rachel borrowed $800 from a bank; she has repaid $2,700 after three years. At what rate of compound interest did the bank lend money to Rachel? (Interest is compounded annually). Hint: A = P (1 + r)n A – amount of money paid after n years P – principal amount r – interest rate n – number of years

OpenStudy (anonymous):

\( \large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~a~\underline{\href{ /study#/updates/527023d6e4b08ad7c0d7b5bc }{guide}}~about~openstudy~here. \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)

OpenStudy (anonymous):

\[ \large 2,700 = 800(1+r)^{3} \]

OpenStudy (anonymous):

\[\large \frac{ 2700 }{ 800 } = (1+r)^3\] \[\LARGE \sqrt[3]{\frac{ 2700 }{ 800 }} =1+r\] \[\LARGE \sqrt[3]{\frac{ 2700 }{ 800 }} -1=r\] r=0.5